OpenStudy (alivejeremy):

A recipe that makes 9 servings calls for cups of juice. How many cups of juice are there per serving? A. B. C. D.

OpenStudy (alivejeremy):

hey can someone help me please

OpenStudy (alivejeremy):

hey ayeedomo

OpenStudy (ayeedomo):

hola

OpenStudy (alivejeremy):

can u help me wit tht question

OpenStudy (alivejeremy):

if u know it

OpenStudy (ayeedomo):

Sorry, I don't know the answer :( I like your icon though :D

OpenStudy (alivejeremy):

lol thts cool

OpenStudy (alivejeremy):

r u a girl

OpenStudy (ayeedomo):

Yup

OpenStudy (alivejeremy):

oh kool

OpenStudy (ayeedomo):

If that's Goku why is his hair blue instead of gold? XD

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @alivejeremy A recipe that makes 9 servings calls for cups of juice. please fill in the missing numbers :) How many cups of juice are there per serving? A. B. C. D. \(\color{#0cbb34}{\text{End of Quote}}\)

Mehek (mehek14):

what are the missing numbers?

OpenStudy (alivejeremy):

tht goku super saiyan god 2

OpenStudy (ayeedomo):

Was this in one of the movies or new season? :/

OpenStudy (alivejeremy):

new season

OpenStudy (alivejeremy):

@Mehek14

Mehek (mehek14):

I believe you need help?

OpenStudy (alivejeremy):

yes i do

Mehek (mehek14):

then please answer what i just asked you

OpenStudy (ayeedomo):

Because I've seen dragon ball, dragon ball z, dragon ball z kai, and gt. Oh that's cool :0 I need to see the new season then!!

OpenStudy (alivejeremy):

@Mehek14 ?

OpenStudy (ayeedomo):

Good luck with answering your question! :)

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 \(\color{#0cbb34}{\text{Originally Posted by}}\) @alivejeremy A recipe that makes 9 servings calls for cups of juice. please fill in the missing numbers :) How many cups of juice are there per serving? A. B. C. D. \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 what are the missing numbers? \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (alivejeremy):

the missing numbers are

OpenStudy (alivejeremy):

@Mehek14 u still there

Mehek (mehek14):

yup

OpenStudy (alivejeremy):

what grade u in

Mehek (mehek14):

9th

OpenStudy (alivejeremy):

oh the missing number is 1/4 with a two beside it

Mehek (mehek14):

\(2\dfrac{1}{4}?\)

OpenStudy (alivejeremy):

yea

Mehek (mehek14):

would it be like this? A recipe that makes 9 servings calls for \(2\dfrac{1}{4}\) cups of juice.

OpenStudy (alivejeremy):

yea

OpenStudy (alivejeremy):

@Mehek14 yea

OpenStudy (alivejeremy):

?

Mehek (mehek14):

so 9c = \(\dfrac{9}{4}\) what are your options?

OpenStudy (alivejeremy):

ok can u help with more

Mehek (mehek14):

if we finish this one...

OpenStudy (alivejeremy):

wym

Mehek (mehek14):

what are the answer choices for this question?

OpenStudy (alivejeremy):

i got it right

Mehek (mehek14):

oh oke

OpenStudy (alivejeremy):

lol

OpenStudy (alivejeremy):

A recipe that makes 6 servings calls for 1/2 cup of butter. Nancy modifies the recipe so that it can serve 8. How many cups of butter does Nancy need?

OpenStudy (alivejeremy):

u get it

OpenStudy (alivejeremy):

@Mehek14

OpenStudy (alivejeremy):

5 years later

Mehek (mehek14):

2/3 should be the answer

OpenStudy (alivejeremy):

i already got it right a year ago

Mehek (mehek14):

oke

OpenStudy (alivejeremy):

lol

OpenStudy (alivejeremy):

A recipe for 1 lb of trail mix calls for 3.2 oz of cashews. How many ounces of cashews are needed to make 4 lb of trail mix? A. 0.8 oz B. 4.2 oz C. 7.2 oz D. 12.8 oz

Mehek (mehek14):

3.2 * 4 = ?

OpenStudy (alivejeremy):

12.4

OpenStudy (alivejeremy):

12.8

Mehek (mehek14):

12.8

OpenStudy (alivejeremy):

Last month, Edwin had $2325.35 that he divided into two bank accounts. He deposited 2/5 into a savings account. He deposited the rest into a checking account. How much money did Edwin deposit into his checking account?

OpenStudy (alivejeremy):

@Mehek14 u still there

Mehek (mehek14):

yes hold on

OpenStudy (alivejeremy):

ok

Mehek (mehek14):

1395.21

OpenStudy (alivejeremy):

Roberto’s monthly pay is $4349.70. Each month he uses 2/5 of his pay for his household expenses and 1/10 for his car payment. He deposits the rest into his savings account. How much does Roberto deposit into his savings account each month?

Mehek (mehek14):

2174.85

OpenStudy (alivejeremy):

thx

Mehek (mehek14):

done?

OpenStudy (alivejeremy):

i had more but u left

OpenStudy (alivejeremy):

but if u still want to help ok

Mehek (mehek14):

oke

OpenStudy (alivejeremy):

hold on

OpenStudy (alivejeremy):

one min

OpenStudy (alivejeremy):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (alivejeremy):

u still there

Mehek (mehek14):

C and D

OpenStudy (alivejeremy):

Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

Mehek (mehek14):

all are correct

OpenStudy (alivejeremy):

u sure

Mehek (mehek14):

yup

OpenStudy (alivejeremy):

A is not

Mehek (mehek14):

0 + (-1) = 0 - 1 = -1

OpenStudy (alivejeremy):

all of them can't be right tho

Mehek (mehek14):

oh wait the question is asking addition

OpenStudy (alivejeremy):

ofc

OpenStudy (alivejeremy):

lol im not dumb

Mehek (mehek14):

whole numbers -> 0+ so no negative numbers

OpenStudy (alivejeremy):

are u asking me tht

Mehek (mehek14):

no im telling you

OpenStudy (alivejeremy):

B,C,D

OpenStudy (alivejeremy):

so its b,c,d

OpenStudy (alivejeremy):

?

Mehek (mehek14):

ye i believe so

OpenStudy (alivejeremy):

sorry i had left

OpenStudy (alivejeremy):

yea i think so to

OpenStudy (alivejeremy):

no its b and d are the only whole numbers

OpenStudy (alivejeremy):

because 0 is not a whole number

OpenStudy (alivejeremy):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. odd natural numbers B. rational numbers C. {0, 1} D. {0, 1, 2}

Mehek (mehek14):

B